# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 5 Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
$\ell_{\bm{\theta}}(x)=\log(\alpha/\sigma)-(x/\sigma)^{-\alpha}-(\alpha+1)\log(x/\sigma),\quad x>0,$ (5.2)
denotes the log density of the Fréchet distribution $P_{(\alpha,\sigma)}$ and where $c>0$ denotes an arbitrary truncation constant. For the case ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$, the rescaled estimation error of $\hat{\bm{\theta}}_{n}^{({\operatorname{mb}})}$ is known to be asymptotically normal (and independent of $c$), that is, under suitable additional assumptions,
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{mb}})}-\alpha_{0}\\ \hat{\sigma}_{n}^{({\operatorname{mb}})}/\sigma_{r_{n}}-1\end{pmatrix}\rightsquigarrow\mathcal{N}_{2}(\bm{\mu},\Sigma^{({\operatorname{mb}})}),$ (5.3)
for some $\bm{\mu}\in\mathbb{R}^{2}$ and some $\Sigma^{({\operatorname{mb}})}\in\mathbb{R}^{2\times 2}$ positive definite, see Bücher and Segers, 2018b ; Bücher and Segers, 2018a . Subsequently, we will extend these results to ${\operatorname{mb}}={\operatorname{cb}}$, and we will derive a consistent bootstrap scheme for the estimation error.
For fixed $k\in\mathbb{N}$, consider the circmax-sample $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$. Independent of the observations, let $\bm{W}_{m(k)}=(W_{m(k),1},\dots,W_{m(k),m(k)})=(W_{1},\dots,W_{m(k)})$ be multinomially distributed with $m(k)$ trials and class probabilities $(m(k)^{-1},\dots,m(k)^{-1})$. As in Section 4, conditional on $\mathcal{M}_{n,r}^{({\operatorname{cb}})}$, the bootstrap sample $\mathcal{M}_{n,r}^{({\operatorname{cb}}),*}$ is obtained by repeating the observations $(M_{r,s}^{({\operatorname{cb}})})_{s\in I_{kr,i}}$ from the $i$th $kr$-block exactly $W_{m(k),i}$-times, for every $i=1,\dots,m(k)$. We are going to show that the independence Fréchet-log-likelihood
${\bm{\theta}}\mapsto\sum_{s=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}}),*}\vee c)=\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}})}\vee c).$
has a unique-maximimizer (with probability converging to one), say $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})$, and that the conditional distribution of the rescaled bootstrap-estimation error, $\sqrt{n/r}(\hat{\alpha}_{n}^{({\operatorname{cb}}),*}-\hat{\alpha}_{n}^{({\operatorname{cb}})},\hat{\sigma}_{n}^{({\operatorname{cb}}),*}/\hat{\sigma}_{n}-1)$, given the observations is close to the distribution of $\sqrt{n/r}(\hat{\alpha}_{n}^{({\operatorname{mb}})}-\alpha_{0},\hat{\sigma}_{n}^{({\operatorname{mb}})}/\sigma_{r_{n}}-1)$ for both ${\operatorname{mb}}={\operatorname{sb}}$ and ${\operatorname{mb}}={\operatorname{cb}}$. We also show how this result can be used to derive valid asymptotic confidence intervals.
A couple of conditions akin to those imposed in Bücher and Segers, 2018b ; Bücher and Segers, 2018a will be needed.
Condition 5.2 (All disjoint block maxima of size $\lfloor r_{n}/2\rfloor$ diverge).
@@
This is Condition 2.2 in Bücher and Segers, 2018a , and guarantees that the probability of the event that all (blocksize $r_{n}$) block maxima in $\mathcal{M}_{n,r}^{({\operatorname{mb}})}$ are larger than $c$ converges to one as well, for ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}},{\operatorname{cb}}\}$. This will guarantee that the Fréchet-log-likelihoods under consideration are well-defined, with probability converging to one. The following condition is Condition 3.4 in Bücher and Segers, 2018b ; recall that $M_{r,1}=\max(X_{1},\dots,X_{r})$.
Condition 5.3 (Moments).
@@
$f_{1}(x)=x^{-\alpha_{0}},\quad f_{2}(x)=x^{-\alpha_{0}}\log x,\quad f_{3}(x)=\log x.$ (5.4)
The first part of this condition on $C_{n}$ is Condition 3.5 in Bücher and Segers, 2018b , and provides control of the bias for those block maxima which are calculated based on $r$ successive observations. For the circmax-sample, the last $r$ observations in a $kr$-block of that sample are not of that form; their contribution to the bias is controlled by the condition on $C_{n,k}$.
@@
$\mathbb{G}_{n,r}^{({\operatorname{cb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{cb}})}-P_{(\alpha_{0},1)}\big{)},\quad\hat{\mathbb{G}}_{n}^{({\operatorname{cb}}),*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),\ast}-\mathbb{P}_{n,r}^{({\operatorname{cb}})})$
@@
$\mathbb{P}_{n,r}^{({\operatorname{cb}})}=\frac{1}{n}\sum_{s=1}^{n}\delta_{(M_{r,s}^{({\operatorname{cb}})}\vee c)/\sigma_{r}},\qquad\hat{\mathbb{P}}_{n,r}^{({\operatorname{cb}}),*}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{(M_{r,s}^{({\operatorname{cb}})}\vee c)/\sigma_{r}};$
@@
Theorem 5.5.
Fix $k\in\mathbb{N}_{\geq 2}$ and $c>0$, and suppose that Conditions 2.2 and 5.15.4 are satisfied. Then, with probability tending to one, there exists a unique maximizer $(\hat{\alpha}_{n}^{({\operatorname{cb}})},\hat{\sigma}_{n}^{({\operatorname{cb}})})$ of the Fréchet log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}})}\vee c)$, and this maximizer satisfies
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{cb}})}-\alpha_{0}\\ \hat{\sigma}_{n}^{({\operatorname{cb}})}/\sigma_{r_{n}}-1\end{pmatrix}=M(\alpha_{0})\mathbb{G}_{n,r}^{({\operatorname{cb}})}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(M(\alpha_{0})(\bm{C}+k^{-1}\bm{C}_{k}),\Sigma^{({\operatorname{sb}})}),$
@@
$M(\alpha_{0})=\frac{6}{\pi^{2}}\begin{pmatrix}\alpha_{0}^{2}&\alpha_{0}(1-\gamma)&-\alpha_{0}^{2}\\ \gamma-1&-(\Gamma^{\prime\prime}(2)+1)/\alpha_{0}&1-\gamma\end{pmatrix},\quad\Sigma^{({\operatorname{sb}})}=\begin{pmatrix}0.4946\alpha_{0}^{2}&-0.3236\\ -0.3236&0.9578\alpha_{0}^{-2}\end{pmatrix}$
@@
Moreover, also with probability tending to one, there exists a unique maximizer $(\hat{\alpha}_{n}^{({\operatorname{cb}}),*},\hat{\sigma}_{n}^{({\operatorname{cb}}),*})$ of ${\bm{\theta}}\mapsto\sum_{s=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{({\operatorname{cb}}),*}\vee c)$, and this maximizer satisfies
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{cb}}),*}-\hat{\alpha}_{n}^{({\operatorname{cb}})}\\ \hat{\sigma}_{n}^{({\operatorname{cb}}),*}/\hat{\sigma}_{n}^{({\operatorname{cb}})}-1\end{pmatrix}=M(\alpha_{0})\hat{\mathbb{G}}_{n,r}^{({\operatorname{cb}}),*}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(0,\Sigma^{({\operatorname{sb}})}).$
Fix ${\operatorname{mb}}\in\{{\operatorname{sb}},{\operatorname{cb}}\}$. If $r_{n}$ is chosen sufficiently large such that, for $j\in\{1,2,3\}$, $C(f_{j})=0$ (${\operatorname{mb}}={\operatorname{sb}}$) or $C(f_{j})=C_{k}(f_{j})=0$ (${\operatorname{mb}}={\operatorname{cb}}$) in Condition 5.4, we have bootstrap consistency in the following sense:
$d_{K}\bigg{[}\mathcal{L}\bigg{(}\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{cb}}),*}-\hat{\alpha}_{n}^{({\operatorname{cb}})}\\ \hat{\sigma}_{n}^{({\operatorname{cb}}),*}/\hat{\sigma}_{n}^{({\operatorname{cb}})}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\bigg{)},\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{({\operatorname{mb}})}-\alpha_{0}\\ \hat{\sigma}_{n}^{({\operatorname{mb}})}/\sigma_{r_{n}}-1\end{pmatrix}\bigg{]}=o_{\mathbb{P}}(1).$
The result in Theorem 5.5 allows for statistical inference on the parameters $\alpha_{0},\sigma_{r}$, for instance in the form of confidence intervals. We provide details on $\sigma_{r}$, using the circular block bootstrap approximation to the sliding block estimator: for $\beta\in(0,1)$, let $q_{\hat{\sigma}_{n}^{*}}(\beta)$ denote the $\beta$-quantile of the conditional distribution of $\hat{\sigma}_{n}^{({\operatorname{cb}}),*}$ given the data, that is, $q_{\hat{\sigma}_{n}^{*}}(\beta)=(F_{\hat{\sigma}_{n}^{*}})^{-1}(\beta)$, where $F_{\hat{\sigma}_{n}^{*}}(x)=\mathbb{P}(\hat{\sigma}_{n}^{({\operatorname{cb}}),*}\leq x\mid\mathcal{X}_{n})$ for $x\in\mathbb{R}$. Note that the quantile may be approximated to an arbitrary precision by repeated bootstrap sampling. Consider the basic bootstrap confidence interval (Davison and Hinkley, , 1997)
$I_{n,\sigma}^{({\operatorname{sb}},{\operatorname{cb}})}(1-\beta)$ $=\big{[}2\hat{\sigma}_{n}^{({\operatorname{sb}})}-q_{\hat{\sigma}_{n}^{*}}(1-\tfrac{\beta}{2}),2\hat{\sigma}_{n}^{({\operatorname{sb}})}-q_{\hat{\sigma}_{n}^{*}}(\tfrac{\beta}{2})\big{]}.$
Corollary 5.6.
Under the conditions of Theorem 5.5 with $r_{n}$ chosen sufficiently large such that, for $j\in\{1,2,3\}$, $C(f_{j})=0$ in Condition 5.4, we have, for any $\beta\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}(\sigma_{r_{n}}\in I_{n,\sigma}^{({\operatorname{sb}},{\operatorname{cb}})}(1-\beta))=1-\beta.$

### v2

@@ Appendix A Application: bootstrapping the pseudo-maximum likelihood estimator for the Fréchet distribution
$\ell_{\bm{\theta}}(x)=\log(\alpha/\sigma)-(x/\sigma)^{-\alpha}-(\alpha+1)\log(x/\sigma),\quad x>0,$ (A.2)
denotes the log density of the Fréchet distribution $P_{(\alpha,\sigma)}$ and where $c>0$ denotes an arbitrary truncation constant. For the case ${\operatorname{m}}\in\{{\operatorname{d}},{\operatorname{s}}\}$, the rescaled estimation error of $\hat{\bm{\theta}}_{n}^{[{\operatorname{m}}]}$ is known to be asymptotically normal (and independent of $c$), that is, under suitable additional assumptions,
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{[{\operatorname{m}}]}-\alpha_{0}\\ \hat{\sigma}_{n}^{[{\operatorname{m}}]}/\sigma_{r_{n}}-1\end{pmatrix}\rightsquigarrow\mathcal{N}_{2}(\bm{\mu},\Sigma^{[{\operatorname{m}}]}),$ (A.3)
for some $\bm{\mu}\in\mathbb{R}^{2}$ and some $\Sigma^{[{\operatorname{m}}]}\in\mathbb{R}^{2\times 2}$ positive definite, see BucSeg18a; BucSeg18-sl. Subsequently, we will extend these results to ${\operatorname{m}}={\operatorname{c}}$, and we will derive a consistent bootstrap scheme for the estimation error.
For fixed $k\in\mathbb{N}$, consider the circmax-sample $\mathcal{M}_{n,r}^{[{\operatorname{c}}]}$. Independent of the observations, let $\bm{W}_{m(k)}=(W_{m(k),1},\dots,W_{m(k),m(k)})=(W_{1},\dots,W_{m(k)})$ be multinomially distributed with $m(k)$ trials and class probabilities $(m(k)^{-1},\dots,m(k)^{-1})$. As in Section 4, conditional on $\mathcal{M}_{n,r}^{[{\operatorname{c}}]}$, the bootstrap sample $\mathcal{M}_{n,r}^{[{\operatorname{c}}],*}$ is obtained by repeating the observations $(M_{r,s}^{[{\operatorname{c}}]})_{s\in I_{kr,i}}$ from the $i$th $kr$-block exactly $W_{m(k),i}$-times, for every $i=1,\dots,m(k)$. We are going to show that the independence Fréchet-log-likelihood
${\bm{\theta}}\mapsto\sum_{s=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{[{\operatorname{c}}],*}\vee c)=\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\ell_{\bm{\theta}}(M_{r,s}^{[{\operatorname{c}}]}\vee c).$
has a unique-maximimizer (with probability converging to one), say $(\hat{\alpha}_{n}^{[{\operatorname{c}}],*},\hat{\sigma}_{n}^{[{\operatorname{c}}],*})$, and that the conditional distribution of the rescaled bootstrap-estimation error, $\sqrt{n/r}(\hat{\alpha}_{n}^{[{\operatorname{c}}],*}-\hat{\alpha}_{n}^{[{\operatorname{c}}]},\hat{\sigma}_{n}^{[{\operatorname{c}}],*}/\hat{\sigma}_{n}-1)$, given the observations is close to the distribution of $\sqrt{n/r}(\hat{\alpha}_{n}^{[{\operatorname{m}}]}-\alpha_{0},\hat{\sigma}_{n}^{[{\operatorname{m}}]}/\sigma_{r_{n}}-1)$ for both ${\operatorname{m}}={\operatorname{s}}$ and ${\operatorname{m}}={\operatorname{c}}$. We also show how this result can be used to derive valid asymptotic confidence intervals.
A couple of conditions akin to those imposed in BucSeg18a; BucSeg18-sl will be needed.
Condition A.2 (All disjoint block maxima of size $\lfloor r_{n}/2\rfloor$ diverge).
@@
This is Condition 2.2 in BucSeg18-sl, and guarantees that the probability of the event that all (blocksize $r_{n}$) block maxima in $\mathcal{M}_{n,r}^{[{\operatorname{m}}]}$ are larger than $c$ converges to one as well, for ${\operatorname{m}}\in\{{\operatorname{d}},{\operatorname{s}},{\operatorname{c}}\}$. This will guarantee that the Fréchet-log-likelihoods under consideration are well-defined, with probability converging to one. The following condition is Condition 3.4 in BucSeg18a; recall that $M_{r,1}=\max(X_{1},\dots,X_{r})$.
Condition A.3 (Moments).
@@
$f_{1}(x)=x^{-\alpha_{0}},\quad f_{2}(x)=x^{-\alpha_{0}}\log x,\quad f_{3}(x)=\log x.$ (A.4)
The first part of this condition on $C_{n}$ is Condition 3.5 in BucSeg18a, and provides control of the bias for those block maxima which are calculated based on $r$ successive observations. For the circmax-sample, the last $r$ observations in a $kr$-block of that sample are not of that form; their contribution to the bias is controlled by the condition on $C_{n,k}$.
@@
$\mathbb{G}_{n,r}^{[{\operatorname{c}}]}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{[{\operatorname{c}}]}-P_{(\alpha_{0},1)}\big{)},\quad\hat{\mathbb{G}}_{n}^{[{\operatorname{c}}],*}=\sqrt{\frac{n}{r}}(\hat{\mathbb{P}}_{n,r}^{[{\operatorname{c}}],\ast}-\mathbb{P}_{n,r}^{[{\operatorname{c}}]})$
@@
$\mathbb{P}_{n,r}^{[{\operatorname{c}}]}=\frac{1}{n}\sum_{s=1}^{n}\delta_{(M_{r,s}^{[{\operatorname{c}}]}\vee c)/\sigma_{r}},\qquad\hat{\mathbb{P}}_{n,r}^{[{\operatorname{c}}],*}=\frac{1}{n}\sum_{i=1}^{m(k)}W_{m(k),i}\sum_{s\in I_{kr,i}}\delta_{(M_{r,s}^{[{\operatorname{c}}]}\vee c)/\sigma_{r}};$
@@
Theorem A.5.
Fix $k\in\mathbb{N}_{\geq 2}$ and $c>0$, and suppose that Conditions 2.2 and A.1A.4 are satisfied. Then, with probability tending to one, there exists a unique maximizer $(\hat{\alpha}_{n}^{[{\operatorname{c}}]},\hat{\sigma}_{n}^{[{\operatorname{c}}]})$ of the Fréchet log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{[{\operatorname{c}}]}\vee c)$, and this maximizer satisfies
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{[{\operatorname{c}}]}-\alpha_{0}\\ \hat{\sigma}_{n}^{[{\operatorname{c}}]}/\sigma_{r_{n}}-1\end{pmatrix}=M(\alpha_{0})\mathbb{G}_{n,r}^{[{\operatorname{c}}]}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(M(\alpha_{0})(\bm{C}+k^{-1}\bm{C}_{k}),\Sigma^{[{\operatorname{s}}]}),$
@@
$M(\alpha_{0})=\frac{6}{\pi^{2}}\begin{pmatrix}\alpha_{0}^{2}&\alpha_{0}(1-\gamma)&-\alpha_{0}^{2}\\ \gamma-1&-(\Gamma^{\prime\prime}(2)+1)/\alpha_{0}&1-\gamma\end{pmatrix},\quad\Sigma^{[{\operatorname{s}}]}=\begin{pmatrix}0.4946\alpha_{0}^{2}&-0.3236\\ -0.3236&0.9578\alpha_{0}^{-2}\end{pmatrix}$
@@
Moreover, also with probability tending to one, there exists a unique maximizer $(\hat{\alpha}_{n}^{[{\operatorname{c}}],*},\hat{\sigma}_{n}^{[{\operatorname{c}}],*})$ of ${\bm{\theta}}\mapsto\sum_{s=1}^{n}\ell_{\bm{\theta}}(M_{r,s}^{[{\operatorname{c}}],*}\vee c)$, and this maximizer satisfies
$\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{[{\operatorname{c}}],*}-\hat{\alpha}_{n}^{[{\operatorname{c}}]}\\ \hat{\sigma}_{n}^{[{\operatorname{c}}],*}/\hat{\sigma}_{n}^{[{\operatorname{c}}]}-1\end{pmatrix}=M(\alpha_{0})\hat{\mathbb{G}}_{n,r}^{[{\operatorname{c}}],*}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow\mathcal{N}_{2}(0,\Sigma^{[{\operatorname{s}}]}).$
Fix ${\operatorname{m}}\in\{{\operatorname{s}},{\operatorname{c}}\}$. If $r_{n}$ is chosen sufficiently large such that, for $j\in\{1,2,3\}$, $C(f_{j})=0$ (${\operatorname{m}}={\operatorname{s}}$) or $C(f_{j})=C_{k}(f_{j})=0$ (${\operatorname{m}}={\operatorname{c}}$) in Condition A.4, we have bootstrap consistency in the following sense:
$d_{K}\bigg{[}\mathcal{L}\bigg{(}\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{[{\operatorname{c}}],*}-\hat{\alpha}_{n}^{[{\operatorname{c}}]}\\ \hat{\sigma}_{n}^{[{\operatorname{c}}],*}/\hat{\sigma}_{n}^{[{\operatorname{c}}]}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\bigg{)},\sqrt{\frac{n}{r}}\begin{pmatrix}\hat{\alpha}_{n}^{[{\operatorname{m}}]}-\alpha_{0}\\ \hat{\sigma}_{n}^{[{\operatorname{m}}]}/\sigma_{r_{n}}-1\end{pmatrix}\bigg{]}=o_{\mathbb{P}}(1).$
The result in Theorem A.5 allows for statistical inference on the parameters $\alpha_{0},\sigma_{r}$, for instance in the form of confidence intervals. We provide details on $\sigma_{r}$, using the circular block bootstrap approximation to the sliding block estimator: for $\beta\in(0,1)$, let $q_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}(\beta)$ denote the $\beta$-quantile of the conditional distribution of $\hat{\sigma}_{n}^{[{\operatorname{c}}],*}$ given the data, that is, $q_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}(\beta)=(F_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}})^{-1}(\beta)$, where $F_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}(x)=\mathbb{P}(\hat{\sigma}_{n}^{[{\operatorname{c}}],*}\leq x\mid\mathcal{X}_{n})$ for $x\in\mathbb{R}$. Note that the quantile may be approximated to an arbitrary precision by repeated bootstrap sampling. Consider the basic bootstrap confidence interval (DavHin97)
$I_{n,\sigma}^{[{\operatorname{s}},{\operatorname{c}}]}(1-\beta)$ $=\big{[}\hat{\sigma}_{n}^{[{\operatorname{s}}]}+\hat{\sigma}_{n}^{[{\operatorname{c}}]}-q_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}(1-\tfrac{\beta}{2}),\hat{\sigma}_{n}^{[{\operatorname{s}}]}+\hat{\sigma}_{n}^{[{\operatorname{c}}]}-q_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}}(\tfrac{\beta}{2})\big{]}.$
Corollary A.6.
Under the conditions of Theorem A.5 with $r_{n}$ chosen sufficiently large such that, for $j\in\{1,2,3\}$, $C(f_{j})=0$ in Condition A.4, we have, for any $\beta\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}(\sigma_{r_{n}}\in I_{n,\sigma}^{[{\operatorname{s}},{\operatorname{c}}]}(1-\beta))=1-\beta.$